OpenStudy (anonymous):

Q:Solve for x: x/(x+2)=(x+1)/(x-3) I feel like im doing it correct but i keep getting x=-1/6 and that is wrong

OpenStudy (anonymous):

did u cross multiply?

OpenStudy (anonymous):

did you "cross multiply"?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol the same

OpenStudy (anonymous):

yea and i got x^2-3x=x^2+3x+1

OpenStudy (anonymous):

i got -2/9

OpenStudy (anonymous):

the last result is wrong

OpenStudy (anonymous):

X^2+3x+1 wrong (x+1)(x+2)= x^2+3x+2

OpenStudy (anonymous):

u just forgot to multiply in the end

OpenStudy (anonymous):

oh wow, haha thank you very much, now i should get -1/3

OpenStudy (anonymous):

k no problem